Under vector addition

α+β+γ=γ+β+α is called the ____ law.


1
Expert's answer
2020-06-15T19:11:20-0400

Commutative


As by commutative property, "a+b = b+c"


similarly, "\alpha + \beta = \beta+ \alpha"


now adding "\gamma" both sides,


"\alpha + \beta + \gamma = (\beta+ \alpha) + \gamma"


"\alpha + \beta + \gamma = \gamma +\beta+ \alpha"